If the numerator and denominator of a fraction are both increased by 5, the fraction becomes 5/6 . If the numerator and denominator are both decreased by one the new fraction becomes2/3 . Find the original fraction.
1
Expert's answer
2020-08-10T18:39:44-0400

Let "\dfrac{a}{b}" be a fraction.


"\dfrac{a+5}{b+5}=\dfrac{5}{6}"

"\dfrac{a-1}{b-1}=\dfrac{2}{3}"

"6(a+5)=5(b+5)""3(a-1)=2(b-1)"

"6a+30=5b+25""3a-3=2b-2"

"6a=5b-5""3a=2b+1"

"2(2b+1)=5b-5""a=\dfrac{1}{3}(2b+1)"

"b=7""a=5"

The original fraction is "\dfrac{5}{7} ."